# Optimizing a Worldwide-Scale Shipper Transportation Planning in a Carmaker Inbound Supply Chain

### v2

@@ 2.2 Commodities and Flows
Most individual shipments fall within a moderate size range of 1 to 4 cubic meters, as shown in figure (3b). While these volumes are manageable, they typically do not fill a transport unit on their own. Consequently, high utilization of containers or trucks depends on consolidating flows across multiple dimensions—such as combining different part types, suppliers, or destinations within the same shipment. This introduces additional decision layers involving timing, compatibility constraints, and routing coordination, especially for multi-stop or multi-leg routes. The current organization of flows results in nearly 500,000 transport units—trucks or shipping containers—being used every six months. Their spatial distribution is shown in Figure 5.
(a) Supplier $\to$ Platform
(b) Platform $\to$ Platform
(c) Platform $\to$ industrial sites
Figure 5: Legs used in current flows
Interestingly, only a small fraction of all potential legs in the network are actively used. This is a direct consequence of the underlying consolidation strategy: flows are channeled through a limited number of high-volume corridors, where cost efficiency can be maximized. In modeling terms, this indicates that leg activation is not static but endogenous—dependent on shipment densities, consolidation opportunities, and transport cost structures. While most of the physical volume moves within Europe, the cost distribution exhibits a contrasting pattern.
Figure 6 illustrates the volume and cost repartition between flows shipped on direct legs and others. The majority of transportation costs stem from intercontinental flows, which are fewer in number but far more expensive per unit. This cost–volume asymmetry has important implications for planning. Intercontinental shipments require long lead times, early capacity reservation, and coordination with deep-sea or multimodal transport services. In contrast, intra-European flows dominate operational complexity due to their higher frequency, network density, and time sensitivity. Intercontinental flows are typically routed through designated global platforms that serve as strategic hubs for mode transitions and consolidation. These hubs enable coordination across continents and help align long-haul shipments with downstream distribution schedules in Europe. Their positioning and function are critical to synchronizing the global and regional components of the supply chain and mitigating the risk of disruptions or bottlenecks.

### v3

@@ A.2 Commodities and Flows
Most individual shipments fall within a moderate size range of 1 to 4 cubic meters, as shown in figure (3b). While these volumes are manageable, they typically do not fill a transport service on their own. Consequently, high utilization of containers or trucks depends on consolidating flows across multiple dimensions—such as combining different part types, suppliers, or destinations within the same shipment. This introduces additional decision layers involving timing, compatibility constraints, and routing coordination, especially for multi-stop or multi-leg routes. The current organization of flows results in nearly 500,000 transport services—trucks or shipping containers—being used every six months. Their spatial distribution is shown in Figure 4. Interestingly, only a small fraction of all potential legs in the network are actively used. This is a direct consequence of the underlying consolidation strategy: flows are channeled through a limited number of high-volume corridors, where cost efficiency can be maximized. In modeling terms, this indicates that leg activation is not static but endogenous—dependent on shipment densities, consolidation opportunities, and transport cost structures. While most of the physical volume moves within Europe, the cost distribution exhibits a contrasting pattern.